OpenStudy (livya15):

I need help refreshing my memory about simplifying a fraction with a fractional exponent in denominator. Will give MEDAL!

OpenStudy (livya15):

\[\frac{ 1 }{ t ^{5/2} }\]

OpenStudy (amonoconnor):

What form are you hoping to get the term into? As one, without any fractions?

OpenStudy (amonoconnor):

You could make it: \[t^{-2.5}\]

OpenStudy (amonoconnor):

That could be a final answer. Do you know how I got there?

OpenStudy (anonymous):

With regards of the exponent, the denominator is the root you are taking, and the numerator will be the exponent under the root. For example, \[x^{1/2} = \sqrt{x}\] and \[x ^{3/2}=\sqrt{x ^{3}}\] Since the term is under 1, that means the exponent must be negative, as in\[x ^{-2}=1/x ^{2}\]